Write the smallest digit and the greatest digit in the blank space of each of the following numbers so that the number formed is divisible by 3:
 4765 __ 2

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Sum of the given digits = 24
Sum of its digits should be divisible by 3 to make the number divisible by 3
Since, 24 is already divisible by 3. Hence, the smallest number that can be replaced is 0 Now, 0 + 3 = 3
3 +3 =6
3 +3 +3 =9
3 + 3 + 3 + 3 = 12
If we put 9, sum of its digits becomes 33.
Since, 33 is divisible by 3.
Therefore the number will be divisible by 3
Hence, the largest number is 9

Thanks